OpenStudy (anonymous):

The expression to find the perimeter of a rectangle is 2(length + width). What is the perimeter, in units, of a rectangle of length fraction 1 over 7 unit and width fraction 1 over 2 unit? fraction 1 and 1 over 7 units fraction 1 and 2 over 7 units fraction 4 over 9 unit fraction 2 over 9 unit

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (phi):

length is 1/7 width is 1/2 can you add those two fractions?

OpenStudy (phi):

\[ \frac{1}{7}+ \frac{1}{2} \] use 14 as the common denominator

OpenStudy (usukidoll):

Perimeter = 2L + 2W let l = 1/7 and w = 1/2 Perimeter =\[2(\frac{1}{7})+2(\frac{1}{2})\]

OpenStudy (anonymous):

ya

OpenStudy (usukidoll):

ok then multiply first.

OpenStudy (anonymous):

multiply what

OpenStudy (usukidoll):

2 x 1/7 + 2 x 1/2

OpenStudy (usukidoll):

what is 2 x 1/7 what is 2 x 1/2

OpenStudy (usukidoll):